can someone please help me with this problem:
Evaluate the function at the given value of the independent variable and simplify.
f(x) = -5x + 8; f(-3)
a. 23
b. -9
c. 3
d. 7

Answer by oscargut(2103)

f(x) = -5x + 8
then
f(-3)=-5(-3)+8=15+8=23
so correct option is a)